This question is about the forces needed to support and to accelerate the train
a)the fully laden transrapid train has a mass of 200 tonnes.What is the force needed to lift it
b)the forward force on the transrapid train is produced by two forces FA and FB.
How the resultant will be horizontal
c)The angle theta is 30 degree and the magnitude of the force FA is 170000N.Calculate the total horizontal Force that FA and FB produce on the train
d)Calculate the acceleration of the train produced by 300000N and suggest a reason why its greater than the .67 ms-2 acceleration of the actual German Transrapid Maglev train whose mass is also 200 tonnes

Grade:

1 Answers

suryakanth AskiitiansExpert-IITB
13 years ago

Dear rameeja,

The force required to lift the train is

F = mg

F = 200*1000*9.81 = 1962000 N

consider angle theta = angle a

Fa+Fb cos(a) = F net

Fb sin(a) = mg = 1962000

Fb* (1/2) = 1962000

Fb = 1962000*2 = 3924000

Fb = 3924000

Fa+Fb cos(a) = F net

F net = 170000 + 3924000*(cos(30)) = 3568283.6844

and acceleration = F/m = 11.89

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Suryakanth –IITB